OpenStudy (anonymous):

HELP!!! :( Which of the following is the simplified version of 1/2 log 3 x - 2 log 3 y + 4 log 3 z ??????

jimthompson5910 (jim_thompson5910):

Are you familiar with any log rules?

OpenStudy (anonymous):

some of them , yes . like the ones you showed me earlier . i just dont understand this problem .

jimthompson5910 (jim_thompson5910):

They just want you to condense this as much as possible

OpenStudy (anonymous):

i'm confused because all of the ansers (except one) have sqaure roots in them .

jimthompson5910 (jim_thompson5910):

The rules we'll use are this... \[\Large y\log_{b}(x) = \log_{b}(x^y)\] \[\Large \log_{b}(x) + \log_{b}(y) = \log_{b}(xy)\] \[\Large \log_{b}(x) - \log_{b}(y) = \log_{b}\left(\frac{x}{y}\right)\]

jimthompson5910 (jim_thompson5910):

We'll also use the idea that \[\Large x^{\frac{1}{2}} = \sqrt{x}\]

OpenStudy (anonymous):

1/2 log3(x) - 2log3(y) + 4log3(z) is that how its supposed to be set up, with the 3 as the base?

jimthompson5910 (jim_thompson5910):

that's one way you can write it

jimthompson5910 (jim_thompson5910):

We have this \[\Large \frac{1}{2}\log_{3}(x)-2\log_{3}(y)+4\log_{3}(z)\] What is the first step?

OpenStudy (anonymous):

i want to say that the first step is taking the square root to get ./x - 2log3(y) + 4log3(z) but i dont think thats right .

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{2}\log_{3}(x)-2\log_{3}(y)+4\log_{3}(z)\] \[\Large \log_{3}\left(x^{\frac{1}{2}}\right)-2\log_{3}(y)+4\log_{3}(z)\] \[\Large \log_{3}\left(\sqrt{x}\right)-2\log_{3}(y)+4\log_{3}(z)\]

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):

either way, log3 is going to be out front, right? so would it be, log3 ./x. can you add -2log3 and +4log3 to get 2log3 ? then combine x & y ?

jimthompson5910 (jim_thompson5910):

Hopefully in step 2, you can see I used the idea that \[\Large y\log_{b}(x) = \log_{b}(x^y)\]

jimthompson5910 (jim_thompson5910):

no you can't add the logs the way you're describing because the stuff in the inside is different

OpenStudy (anonymous):

ohh. so... i'm lost again :/ math isnt my strong point :(

jimthompson5910 (jim_thompson5910):

sry made a typo, but I fixed it

jimthompson5910 (jim_thompson5910):

oh wait...one sec

OpenStudy (anonymous):

okay. hmmm... is it log3 ./xz^4 over y^2 ?

jimthompson5910 (jim_thompson5910):

sry, i write it out,but then I realized I made a typo (and I can't go back and fix it). So i have to erase and rewrite \[\Large \log_{3}\left(\sqrt{x}\right)-2\log_{3}(y)+4\log_{3}(z)\] \[\Large \log_{3}\left(\sqrt{x}\right)+4\log_{3}(z)-2\log_{3}(y)\] \[\Large \log_{3}\left(\sqrt{x}\right)+\log_{3}(z^4)-2\log_{3}(y)\] \[\Large \log_{3}\left(\sqrt{x}*z^4\right)-2\log_{3}(y)\] \[\Large \log_{3}\left(z^4\sqrt{x}\right)-2\log_{3}(y)\]

OpenStudy (anonymous):

YAY! so i got it right??? :D

jimthompson5910 (jim_thompson5910):

yes, keep going to get... \[\Large \log_{3}\left(z^4\sqrt{x}\right)-2\log_{3}(y)\] \[\Large \log_{3}\left(z^4\sqrt{x}\right)-\log_{3}\left(y^2\right)\] \[\Large \log_{3}\left(\frac{z^4\sqrt{x}}{y^2}\right)\]

OpenStudy (anonymous):

sweet!! thank you!!!

jimthompson5910 (jim_thompson5910):

sure thing